### v1

@@ C.3 Shallow water Equations
C.4 Zero-shot super-resolution results
@@
Table 3: Zero-shot super-resolution results for Darcy flow and the spherical shallow water problems. The validation error is reported in terms of relative $L^{2}$-error at double the training resolution. Autoregressive rollouts are super-resolved in the sense that the rollout is performed at the high resolution.
Model Parameters Relative $L^{2}$-Error. (super-res.)
# Layers # Modes Embedding # Parameters 1 step 5 steps
@@
FNO 4 20 41 $2.715\cdot 10^{6}$ $8.646\cdot 10^{-2}$
FNO + diff. kernel (ours) 4 20 40 $2.599\cdot 10^{6}$ $\mathbf{7.774\cdot 10^{-2}}$
@@
Spherical U-Net (with local integral kernel) 17 - 32 $1.639\cdot 10^{6}$ $3.386\cdot 10^{-3}$ $1.244\cdot 10^{-2}$
SFNO 4 128 32 $1.066\cdot 10^{6}$ $3.830\cdot 10^{-3}$ $1.427\cdot 10^{-2}$
SFNO + local integral kernel (ours) 4 128 31 $1.019\cdot 10^{6}$ $\mathbf{3.232\cdot 10^{-3}}$ $\mathbf{7.450\cdot 10^{-3}}$
@@
In the Darcy flow setting, we use the same setup and dataset as described in Section 4.1. We train two models on data sampled at a $256\times 256$ regular grid and evaluate on a $512\times 512$ regular grid (2x super-resolution and 4x the number of points). For the shallow water equations, a similar approach is taken. We take the models in Section 4.3 which were trained at a resolution of $256\times 512$ and apply them to data generated at a resolution of $512\times 1024$.
Table 3 lists the results of our super-resolution experiment with our proposed models. In the Darcy setting, we compare the baseline FNO to the FNO augmented with the differential kernel. On the sphere, we make use of the fact that the spherical U-Net presented in Ocampo et al. (2022) is already a neural operator due to its discretization-independence. As such, we can use it alongside the SFNO as a baseline for zero-shot super-resolution on the sphere.

### v2

@@ C.4 Shallow water Equations
C.5 Flow past a cylinder
To demonstrate the capability of dealing with unstructured representations, we use the dataset provided by Rahman et al. (2024) to train a model for predicting velocity and pressure fields for a Navier-Stokes problem in a channel with a suspended cylinder and attached membrane (see Figure 5). We use the dataset with a viscosity of $\mu=1.0$, which corresponds to a Reynolds number of $\text{Re}=2000$. To deal with the unstructured data, we employ a local integral convolution layer in both the encoder and decoder with a cutoff radius of $r_{\text{cutoff}}=0.052$ and 5 piecewise-linear basis functions. This layer transforms the unstructured data to a regularly spaced grid of $48\times 192$ in the internal representation. The processor part of the architecture then consists of 4 FNO blocks with local integral kernels, which have the same filter basis as the convolutions used in the encoder/decoder. The embedding dimension is fixed at 16 and all Fourier modes are kept in the internal representation. Overall, this leads to a parameter count of $9.6\cdot 10^{6}$. The architecture is trained on a dataset of 250 samples for 50 epochs using the Adam optimizer and a learning rate of $2\cdot 10^{-3}$. The results alongside baselines from Rahman et al. (2024) are reported in Table 3.
C.6 Zero-shot super-resolution results
@@
Table 5: Zero-shot super-resolution results for Darcy flow and the spherical shallow water problems. The validation error is reported in terms of relative $L^{2}$-error at various resolutions. Autoregressive rollouts are super-resolved in the sense that the rollout is performed at the high resolution.
Model Parameters Relative $L^{2}$-Error
# Layers # Modes Embedding # Parameters resolution (relative to training resolution)
1/2×1/2$\times$1 / 2 × 1×1$\times$1 × 2×2$\times$2 × 4×4$\times$4 ×
@@
FNO 4 20 41 $2.715\cdot 10^{6}$ $1.475\cdot 10^{-1}$ $5.867\cdot 10^{-2}$ $8.646\cdot 10^{-2}$ $7.731\cdot 10^{-2}$
FNO + diff. (ours) 4 20 40 $2.599\cdot 10^{6}$ $\mathbf{1.174\cdot 10^{-1}}$ $\mathbf{5.851\cdot 10^{-2}}$ $\mathbf{7.774\cdot 10^{-2}}$ $\mathbf{6.681\cdot 10^{-2}}$
@@
Spherical U-Net 17 - 32 $1.639\cdot 10^{6}$ $5.586\cdot 10^{-3}$ $6.160\cdot 10^{-4}$ $3.386\cdot 10^{-3}$ $4.102\cdot 10^{-3}$
SFNO 4 128 32 $1.066\cdot 10^{6}$ $1.342\cdot 10^{-3}$ $9.220\cdot 10^{-4}$ $3.830\cdot 10^{-3}$ $4.419\cdot 10^{-3}$
SFNO + loc. int. (ours) 4 128 31 $1.019\cdot 10^{6}$ $\mathbf{8.673\cdot 10^{-4}}$ $\mathbf{2.624\cdot 10^{-4}}$ $\mathbf{3.341\cdot 10^{-3}}$ $\mathbf{4.097\cdot 10^{-3}}$
@@
In the Darcy flow setting, we use the same setup and dataset as described in Section 4.1. We train two models on data sampled at a $256\times 256$ regular grid and evaluate on a $128\times 128$ regular grid ($\frac{1}{2}$x resolution and $\frac{1}{4}$x the number of points), $512\times 512$ grid (2x super-resolution), and $1024\times 1024$ grid (4x super-resolution). For the shallow water equations, a similar approach is taken. We take the models in Section 4.4 which were trained at a resolution of $256\times 512$ and apply them to data generated at a resolution of $128\times 256$, $512\times 1024$, and $1024\times 2048$.
Table 5 provides the details of the models we used for these experiments as well as the results of our super-resolution experiments with our proposed models. In the Darcy setting, we compare the baseline FNO to the FNO augmented with the differential kernel. On the sphere, we make use of the fact that the spherical U-Net presented in (Ocampo et al., 2022) is already a neural operator due to its discretization-independence. As such, we can use it alongside the SFNO as a baseline for zero-shot super-resolution on the sphere.
@@
C.7 Computational efficiency of differential and local integral kernels
On equidistant grids, our differential and local integral kernels can be implemented as standard convolutional kernels, which are heavily optimized for GPU performance. The computational complexity of our differential kernels is linear in the number of grid points since the size of the convolutional kernel remains constant regardless of resolution. In contrast, the size of the discretized local integral kernels does increase with input resolution. While this may be expensive for some high-resolution datasets, in our experiments, we found that the heavily-optimized convolutions in deep learning libraries can help reduce this computational burden. As a consequence, local integral kernels are more efficient than GNO (Li et al., 2020b) and other graph-based approaches on equidistant grids.
Moreover, the local integral kernels can also be applied to irregular grids, where they are implemented as sparse matrix multiplications. This tends to outperform GNOs (Li et al., 2020b), which need to evaluate a neural network on the graph. The complexity of the local integral layer is linear with a constant that depends on the sparsity of the matrix. This, in turn, depends on the support of the local kernel; choosing a fixed support will scale quadratically with the resolution, whereas a support that matches the resolution will scale linearly in practice.